OpenStudy (anonymous):

for what values of x, y and z is 2-2x=y -3z+4x=0 9z-6=-4y true?

OpenStudy (anonymous):

Tell me what to do.

OpenStudy (anonymous):

can you solve this using substitution? can you plug 2-2x=y into 9z-6=-4y so it equals 9z-6=-4(2-2x)?

OpenStudy (anonymous):

Multiply first equation by 2 and then add equation 1 with equation 2.. you will get another equation having y and z as two variables in it..

OpenStudy (anonymous):

Yes you can do that.. but it will go longer..

OpenStudy (anonymous):

Or you wanna solve this by Substitution Method??

OpenStudy (anonymous):

I rather do it the quickest way :) the actual problem did not specify which was it wants us to solve it

OpenStudy (anonymous):

Then you must use Elimination Method... Multiply first equation by 2 and tell me what did you get??

OpenStudy (anonymous):

4-4x=2y?

OpenStudy (anonymous):

Yep now add this equation with second one and tell me what are you having now??

OpenStudy (anonymous):

3z+4=2y?

OpenStudy (anonymous):

It should come -3z there Right ??

OpenStudy (anonymous):

oh yea! sorry -3z+4=2y

OpenStudy (anonymous):

Rearrange it so that it must look nice to me... Can you do that?

OpenStudy (anonymous):

4-3z=2y? or 2y+3z=4?

OpenStudy (anonymous):

Second one is looking very nice to me..

OpenStudy (anonymous):

Variables on one side and constants on other side..

OpenStudy (anonymous):

Now rearrange the third equation also.. (of your question)..

OpenStudy (anonymous):

9z+4y=6?

OpenStudy (anonymous):

or : 4y + 9z = 6..

OpenStudy (anonymous):

Now you have : 2y + 3z = 4 4y + 9z = 6 Now multiply first equation with 2..

OpenStudy (anonymous):

4y+6z=8?

OpenStudy (anonymous):

Now subtract it with 2nd equation and find z..

OpenStudy (anonymous):

4y + 6z = 8 4y + 9z = 6 Subtract them..

OpenStudy (anonymous):

-2/3?

OpenStudy (anonymous):

Yep.. Now find y carefully by putting z in any one equation..

OpenStudy (anonymous):

Getting ???

OpenStudy (anonymous):

y=1?

OpenStudy (anonymous):

No check it one more time, you are going really well..

OpenStudy (anonymous):

Let us put it in : 2y + 3z = 4 Here put z = -2/3..

OpenStudy (anonymous):

y=3?

OpenStudy (anonymous):

Yep now you got it right.. Can you find from first equation now?? (of your question)..

OpenStudy (anonymous):

*can you find x..

OpenStudy (anonymous):

x=-1/2?

OpenStudy (anonymous):

Yep great job @kkvoo You are a good learner...

OpenStudy (anonymous):

Hooray! Thank you :D

OpenStudy (anonymous):

You are welcome dear...